Hello,
Please help me with the following problem:
1. Solve the equation by the method of your choice.
5x^2-18x=8
The solution set is_________ (Simplify your answer. Type integers or fractions. Use a comma to separate answers.)

Answer by scott8148(6628)

5x^2 - 18x - 8 = 0

factoring ___ (5x + 2) (x - 4) = 0

5x + 2 = 0 ___ x = -2/5

x - 4 = 0 ___ x = 4